40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 2166     Entry time: Fri Apr 6 15:57:30 2018
Author: awade 
Type: DailyProgress 
Category: Other 
Subject: Tuning up alignement into RefCavs 

Since hard clamping down the vaccan, the alignment into the cavities has been suboptimal. 

South path was down to 45% fringe vis and north was down to 14%.  As we are using the trans PDs as reference for autolocker thresholds this is causing the FSS to unlock periodically.

I walked the south periscope so that the reflected DC voltage was V_min, Vmax = (0.292 V, 1.52  V) => Vis = 67%

Similarly the north was improved V_min, Vmax = (0.376V, 1.76V) => Vis 65%.

ELOG V3.1.3-